OpenStudy (anonymous):

16. Consider the line that passes through the points (–5, 9) and (–2, 0). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y – 9 = –(x + 5) y – 9 = –3(x + 5) (4 points) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (anonymous):

the second answer is the right one

OpenStudy (anonymous):

U sure

OpenStudy (anonymous):

yes. first you find the slope of the line: y2-y1 all over x2 - x1 which is: 9-0/ (-5 - (-2)) =9/-3 =-3

OpenStudy (anonymous):

then you substitute the slope and one of the points into point slope formula

OpenStudy (anonymous):

which is equation #2

OpenStudy (anonymous):

Y- 9 =-3(x+5)